OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP!! Just need to finish just 1 and I'm done!!

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

ummmmmm

OpenStudy (anonymous):

Calculate the area of triangle WXY with altitude YZ, given W (2,-1), X (6,3), Y (7,0) and Z (5,2) 8 square units 9.4 square units 7.7, square units 12 square units

OpenStudy (anonymous):

@mumustar78612

OpenStudy (campbell_st):

well I'd say you need the distance formula to find the base length and altitude the base would be WX, so find the distance from (2, -1) tp (6,3) and the altitude YZ is the distance from (7,0) to (5,2) then use the normal area of a triangle A =1/2bh

OpenStudy (anonymous):

I got 7.7 idk if that's right

OpenStudy (campbell_st):

I think you need to check your answer

OpenStudy (anonymous):

Is it 9.4?

OpenStudy (campbell_st):

\[WX = \sqrt{(6 -2)^2 + (3 - (-1))^2}\] \[YZ = \sqrt{(7 - 5)^2 + (0 - 2)^2}\] find those 2 values..

OpenStudy (anonymous):

I got 4

OpenStudy (anonymous):

Well 32/8

OpenStudy (anonymous):

I calculated everything

OpenStudy (campbell_st):

\[WX = \sqrt{32} = 4\sqrt{2}~~~and~~~YZ = \sqrt{8} = 2\sqrt{2}\] then the area is \[A = \frac{1}{2} \times WX \times YZ\] which is \[A = \frac{1}{2} \times 4 \sqrt{2} \times 2 \sqrt{2}\]

OpenStudy (anonymous):

12??

OpenStudy (campbell_st):

so you have \[A = \frac{1}{2} \times 4 \sqrt{2} \times 2\sqrt{2} = 2 \times 2 \times \sqrt{2} \times \sqrt{2}\] and you need to know \[\sqrt{a} \times \sqrt{a} = a\] hope that helps